OpenStudy (anonymous):

x(5x-3)=26 Solve this, is this prime? If not, how did you get the answer? Thanks.

OpenStudy (anonymous):

It's not prime

OpenStudy (anonymous):

darn

OpenStudy (anonymous):

5x^2-3x-26=0 Use the quadratic equation

OpenStudy (anonymous):

5x^2 -3x -26=0 5x^2-13x+10x-26=0 (x+2)(5x-13)=0

OpenStudy (anonymous):

Oh! I see what i did wrong now, thanks :D

OpenStudy (anonymous):

Or you can factor by grouping like notsobright did.

OpenStudy (anonymous):

Thats what I usually do, the grouping thing, I just didnt do something correctly :P